# Advances in Embodied Navigation Using Large Language Models: A Survey

### v4

@@ 4 Challenges and Future Directions
5 Conclusion
This paper delves into the rapidly evolving field of Embodied Intelligence, focusing on the premise that intelligence emanates from the interaction between an agent and its environment rather than being a purely internal, abstract construct. The paper scrutinizes 14 works in embodied navigation, delineating their variances and commonalities. Furthermore, it offers an in-depth analysis of how these works catalyze contemporary advancements in research tasks in embodied AI, shedding light on their underlying design principles and methodologies. Experimental outcomes indicate significant success rate improvements across various work evaluations.
The paper also outlines challenges and limitations inherent to LLMs in the realm of embodied intelligence. These include the absence of a direct linkage between LLMs and the physical world, the necessity for extensive training data, and the complexity of understanding and generating natural language in a diverse array of contexts. Notwithstanding these impediments, several promising research trajectories exist that could propel advancements in embodied intelligence. These encompass the formulation of more robust and adaptable language models, the investigation of innovative training methodologies and architectures, the establishment of standardized works and evaluation metrics, and the critical need for cross-disciplinary collaboration among researchers in AI, robotics, and social sciences.

### v5

@@ IV Challenges and Future Directions
V Conclusion
This paper delves into the rapidly evolving field of Embodied Intelligence, focusing on the premise that intelligence emanates from the interaction between an agent and its environment rather than being a purely internal, abstract construct. The paper scrutinizes several works in embodied navigation, delineating their variances and commonalities. Furthermore, it offers an in-depth analysis of how these works catalyze contemporary advancements in research tasks in embodied AI, shedding light on their underlying design principles and methodologies. The paper also outlines challenges and limitations inherent to LLMs in the realm of embodied intelligence. These include the absence of a direct linkage between LLMs and the physical world, the necessity for extensive training data, and the complexity of understanding and generating natural language in a diverse array of contexts. Notwithstanding these impediments, several promising research trajectories exist that could propel advancements in embodied intelligence. These encompass the formulation of more robust and adaptable language models, the investigation of innovative training methodologies and architectures, the establishment of standardized works and evaluation metrics, and the critical need for cross-disciplinary collaboration among researchers in AI, robotics, and social sciences.